# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v1

@@ 7. Research Cases
To intuitively demonstrate the deployment potential and challenges faced by NSVAD in real-world applications, we selected three of our research cases to apply NSVAD in modern industry (Liu et al., 2024b), smart cities (Liu et al., 2023c), and large-scale systems.
@@ 7.1. Appearance-Motion Prototype Network for Smart Industry
NSVAD can gather global spatial-temporal information in open spaces and detect various anomalous events in industrial settings to maintain production process safety and stability. Compared to traditional monitoring devices, video camera data offers stable quality and high information density, making NSVAD the preferred solution for Automatic Industrial Anomaly Detection Systems (AIADS). However, existing VAD research lacks in-depth exploration of industrial scenarios, and proposed methods often work only on outdoor public spaces, struggling to cope with the high-temporal dynamics and complex spatial-temporal interactions in manufacturing sites. On the one hand, the scarcity of industrial video datasets limits the development of AIADS. On the other hand, the diversity of industrial events and the complexity of patterns in such dynamic environments make developing efficient industrial NSVAD systems capable of effectively representing regular patterns and limitedly generalizing anomaly events a highly challenging task. To address this, inspired by the categorization of UVAD anomalies in Section 2.2 and motivated by the multi-stream approach introduced in Section 3.1., we propose in (Liu et al., 2024b) the first NSVAD model for AIADS, named the Appearance-Motion Prototype Network (AMP-Net), as shown in Fig. 10. Quantitative experiments on real-world industrial datasets demonstrate that AMP-Net can handle dynamic environments while maintaining limited generalization to various industrial anomalies.
As the first NSVAD case tailored for industrial scenarios, AMP-Net demonstrates the feasibility of using video for automated monitoring of production processes and the application prospects of NSVAD in the smart industry. The implementation of this model would not have been possible without previous explorations, particularly in multi-stream approaches (Liu et al., 2022c) and memory networks (Gong et al., 2019; Park et al., 2020), which also inspire future researchers to seek inspiration from existing solutions tailored to similar tasks when designing NSVAD systems for specific applications, as explained in Section 3.1 of Supplement. Although the model achieves excellent performance on self-collected industrial datasets, it fails to consider the deployment cost on terminal devices and the arbitrary data shifts in real-world environments, indicating that there is still a distance to go before it can be truly applied in modern factories. Additionally, quantitatively analyzing the model’s detection capabilities for different industrial anomaly events to develop interpretable NSVAD systems and improve the model’s detection capabilities for highly hazardous anomalies is also a realistic requirement worth exploring.
@@ 7.2. Causal Video Normality Learning Towards Smart Cities
Smart cities are reshaping human lifestyles, relying on data sensing and understanding technicals driven by AI and next-generation communication. Among the vast and multimodal data generated by typical smart city applications such as modern transportation, smart homes, and public services, videos often dominate data volume and information density, making automated video processing a fundamental task in smart city research. NSVAD, tailored for smart cities, holds the promise of providing new solutions for traffic accident detection, elderly state monitoring, and public safety protection. However, challenges such as label-independent data shifts and subtle anomalies brought about by the diversity of events in such complex scenarios need to be addressed. Despite efforts by UVAD methods (Liu et al., 2018b; Gong et al., 2019; Chang et al., 2020), including the AMP-Net (Liu et al., 2024b), to balance the representation and generalization capabilities, deep learning-based models have struggled to establish simple mapping relationships between regular events and labels, lacking exploration of the causal relationships in video regularity. Our case studies in (Liu et al., 2023c) indicate that deep representation learning (DRL) struggles to learn the essential differences between normal and abnormal events and fails to resist arbitrary biases in smart cities, leading to significant performance degradation when handling complex regular events and minor anomalies.
To this end, we introduced Causal Representation Learning (CRL) into NSVAD, proposing Causal Video Normality Learning (CVNL) to explore potential causal relationships (Liu et al., 2023c). Compared to the shallow description of video normality based on statistical dependency established by DRL, CVNL attempts to discern anomalies based on the inherent causal consistency of regular events, enabling NSVAD to generalize reasonably well to diverse regular events with label-independent biases while remaining sensitive to anomalies. Extensive experiments on large-scale cross-scene datasets demonstrate that CVNL can meet the detection requirements in complex environments like smart cities. Even when only a few regular events are available, it can maintain stable performance. The simplified framework of CVNL is outlined in Fig. 11, and more information is provided in the paper (Liu et al., 2023c) and Section 3.2 of Supplement.
@@ 7.3. End-to-Cloud Collaborative Framework for NSVAD
Inspired by NSAI (Song et al., 2022), we delineate the research on deployable NSVAD systems for real-world, large-scale IoT and mobile internet applications into a hierarchical structure as shown in Fig. 12, comprising hardware layer, system layer, algorithm layer, and application layer. Existing NSVAD works mainly focus on the algorithm and application, emphasizing model development for specific scenarios and aiming to design customized solutions such as unsupervised crowd anomaly discrimination, weakly supervised crime identification, and multimodal violence detection. Due to the limited computational resources of terminal cameras, such models typically rely on edge devices of local monitoring systems to perform data analysis and anomaly detection, resulting in limited spatial scale and temporal length of detectable anomalies. Taking traffic management in smart cities as an example, constructing NSVAD systems for millions of traffic cameras across the entire metropolitan area is necessary for detecting traffic congestion affecting overall road conditions and long-term abnormal traffic flows. However, existing methods can only achieve local short-term detection at the intersection level or even on single terminal devices. Although the development of mobile communication technology enables smart terminals such as mobile phones and tablets to utilize cloud servers for device-interconnected large-scale model learning and global detection, existing cloud-based mobile applications typically upload raw data, which is feasible for NSVAD. On the one hand, the information density of video data and the data set that can be collected per unit of time are higher than other modalities, leading to a significant increase in communication overhead for mobile devices. On the other hand, anomaly detection tasks often involve sensitive information, and directly transmitting such data to the cloud may raise concerns about user privacy among the public. To address this, to balance the communication overhead between terminal devices and cloud servers and eliminate data security risks, we design a configurable end-to-cloud collaborative framework.
Several end devices and a cloud server collaborate to accomplish various types of anomaly event detection, including local, global, instantaneous, and long-term, for large-scale terminal systems. Specifically, monitoring cameras or mobile devices on the end-side can serve as computing relays for cloud servers to perform preprocessing and feature extraction tasks. Thus, end devices tailored for specific scenarios or users only need to transmit data features or intermediate results to the cloud, significantly reducing sensitivity compared to raw visual information. The cloud server aggregates features and results from various edn devices and constructs large models (Liu et al., 2023b) with global and long-term detection capabilities, unloading models to local devices or directly reporting results to the end according to scene requirements. Our end-cloud collaborative NSVAD architecture is flexible and configurable, supporting an indefinite number of devices to join, and the execution party of each task can freely choose between end and cloud according to actual needs (such as real-time performance and complexity). Thus, this architecture can leverage the natural advantages of devices close to raw data, thereby reducing latency and communication overhead, while also utilizing cloud computing power to build high-performance large models without raising concerns about data privacy.
@@
The core modules of the ECC-NSVAD system include a deployment platform, data pipeline, and computing containers. Computing containers aim to provide a cross-platform and lightweight execution environment for various NSVAD models, supporting continuous updates of models. The deployment platform directly targets terminal devices such as monitoring cameras and smartphones, deploying large NSVAD models learned from the cloud locally, similar to existing end-cloud collaborative architectures used for e-commerce recommendation services. The data pipeline is based on the Flink stream processing framework, which directly uploads all raw data collected by devices to the cloud, leading to high redundancy, large communication overhead, and high privacy sharing. To address this, we propose a lightweight data pipeline referring to Alibaba’s Walle framework, which optimizes resource overhead through end-cloud collaboration at the data level. The real-time channel adopts long-link communication and optimizes SSL protocol to reduce latency. In practical applications, ECC-NSVAD utilizes lightweight models to execute instantaneous local anomaly behavior detection on end devices and transfers features of low-confidence segments to the cloud for detecting potential complex anomalies. Case studies show that compared to traditional data center solutions (where all videos are directly transmitted to the cloud for central model processing), our end-cloud collaborative framework reduces cloud workload by 78%, increases detection accuracy (frame-level AUC) by 12%, and reduces false positive rates by 46% in complex dynamic environments.
@@ 8.1. Research Challenges
In Sections 3-5, we have briefly outlined scientific problems addressed by various algorithms when presenting their motivation and implementation. To further help readers grasp the key challenges in NSVAD research, we detailed the unresolved and impending problems from the perspectives of data, labels, models, and systems. Unlike previous works focusing solely on algorithm design, we also discussed the cost bottlenecks that NSVAD deployments may encounter in real-world applications, aiming to guide researchers from different fields.
@@ 8.1.1. Data
Due to variations in scenes, camera view, and device configurations, real-world videos often exhibit label-independent domain shifts, where videos with the same semantic labels manifest slight differences in spatial-temporal patterns. Although such variations are typically easily understood by humans, they may serve as the basis for discerning anomalies by NSVAD models, leading to high false positive rates. Existing methods often validate models on datasets collected from a single scene with the same device and view to circumvent this issue. Although the UCSD dataset (Li et al., 2013) captures videos from two perspectives, parallel to and perpendicular to intersections, it is divided into two datasets for learning view-specific NSVAD models. Additionally, the ShanghaiTech dataset (Liu et al., 2018b) comprises 13 scenes, yet existing methods often overlook scene differences, treating it as single-scene, resulting in significantly lower performance compared to simple datasets like UCSD Ped2 (Li et al., 2013) and CUHK Avenue (Lu et al., 2013). In the real world, variations in scenes and views, as well as disparities between surveillance cameras and smartphones, are inevitable, making it unrealistic to learn specialized models for each specific setup. Given the rarity of anomalies, recent researchers have proposed using virtual data engines to simulate anomaly events to construct rich positive sample sets. Additionally, the XD-Violence dataset (Wu et al., 2020) introduces movie and game scenes. While such data enables the training of large-scale models, they often differ from real anomalies, rendering trained models ineffective in real-world applications. Therefore, bridging the domain shift between virtual anomalies and the real world is crucial for deployable NSVAD models. Despite becoming a hot topic in anomaly detection and multimodal understanding communities, multimodal NSVAD is still limited to the fusion of RGB images and synchronized audio. Existing methods focus solely on traditional image, audio, and optical flow modalities, unable to extract anomaly clues from new modalities like language and subtitle text, limiting NSVAD applications in streaming media and video broadcasting.
@@ 8.1.2. Label
Despite the repeated emphasis in the anomaly detection community on the rarity and diversity of anomaly samples, mainstream anomaly detection models often train in an unsupervised manner using only normal samples. However, as stated in Section 2.1, unsupervised methods actually require all training sample labels to be 0. Therefore, challenges in the label aspect include the impact of noisy labels on model performance and the annotation cost of large-scale datasets. Specifically, UVAD relies on additional data filtering to avoid mixing anomaly samples in the training set, making it unable to perform online learning directly from raw video streams like FuVAD and continuously improve model detection capabilities in real-world environments. Another type, WsVAD, relies on video-level labels to avoid high labor costs. Future work should focus on improving the stability of FuVAD and its performance in complex environments or developing novel hybrid supervised models to eliminate the influence of label noise.
@@ 8.1.3. Model
Unsupervised NSVAD benefits from efficient representation of spatial-temporal features by video self-supervised learning models but may also miss anomalies due to over-generalization. Therefore, the key of unsupervised methods lies in balancing representation and generalization to simultaneously reduce false positives and false negatives. Cases in Fig. 10 and 11 demonstrate the feasibility of memory networks and causal representation learning to achieve this goal, but performance differences on complex datasets indicate that improving the adaptability of UVAD in real environments remains a challenge. Additionally, WsVAD can only detect predefined specific anomaly events, violating the open-world assumption. This not only limits the application scenarios of the model but also hinders the collaborative development of NSVAD technology, leading existing reviews to typically not consider it within the scope of VAD research. Although existing research indicates that the detection results of WsVAD are more reliable compared to UVAD and FuVAD, its core ranking loss is not convincingly accepted. To further promote VAD technology in smart surveillance systems and mobile internet applications, it is necessary to develop interpretable models and rely on NSVAD systems with data encryption capabilities.
@@ 8.1.4. System
Existing NSVAD research focuses solely on improving performance on existing datasets while ignoring a series of challenges in model deployment, such as the compatibility of heterogeneous devices and resource costs. Specifically, NSVAD for different businesses (such as surveillance video analysis, movie content monitoring, and non-real video identification) needs to deal with different task settings and data modalities. However, existing work mainly focuses on anomaly detection in surveillance videos while neglecting new applications like short videos and live streaming. Additionally, terminal devices, especially mobile phones, often have limited storage, computing, and communication resources, so NSVAD model design needs to move towards lightweight and shift from algorithm to system to achieve synchronous optimization of detection performance and cost.

### v2

@@ 6.1.2. Cost-oriented metrics
6.2. System-level Evaluation
While conventional metrics focus on the detection accuracy and cost of individual models, evaluating the performance of NSVAD in real-world deployments requires a more holistic, system-level perspective. Beyond model-specific metrics, the entire system’s effectiveness depends on various factors such as latency, communication cost, bandwidth efficiency, data security, user privacy, and system robustness. These metrics collectively assess the performance and feasibility of deploying NSVAD in large-scale, distributed, and resource-constrained environments. To this end, we categorize system-level metrics into three primary groups: efficiency, privacy, and robustness.
6.2.1. Efficiency-oriented metrics
Efficiency-oriented system metrics assess the efficiency and scalability of the NSVAD system in a distributed setting. One key metric is latency, which measures the end-to-end delay from the moment video data is captured to the final detection output. For real-time anomaly detection, minimizing latency is critical, particularly in scenarios such as public safety monitoring or autonomous driving, where any detection delay could result in catastrophic outcomes. Latency can be broken down into communication latency, processing latency, and system response time, each reflecting different aspects of delay within the system. Reducing these latencies often requires optimizing the placement of inference tasks across edge and cloud nodes. Communication cost evaluates the data transmission overhead between distributed nodes, especially in edge-cloud architectures. Given the high bandwidth demands of video data, optimizing communication efficiency becomes crucial for deploying NSVAD at scale. Common metrics used here include the total data transmitted (measured in megabytes or gigabytes) and the number of communication rounds required for model updates in federated learning-based systems. To address these challenges, techniques such as video compression, parameter pruning, and model quantization are often employed to reduce communication overhead. Another essential metric is bandwidth utilization, which measures the efficiency of the network resources. High bandwidth usage may congest the network, causing delays and performance degradation, particularly in multi-client or large-scale environments. Methods such as asynchronous communication and bandwidth allocation prioritization can be leveraged to optimize utilization without sacrificing detection performance.
6.2.2. Privacy-oriented metrics
In NSVAD systems, maintaining data security and user privacy is paramount, particularly in applications involving sensitive environments such as healthcare (Li et al., 2024c) or public surveillance. Data security is typically evaluated using metrics such as encryption overhead, which measures the computational cost introduced by encryption techniques, and key management efficiency, which assesses how well the system handles the distribution and renewal of cryptographic keys in a large-scale deployment. Robust encryption algorithms, such as AES or homomorphic encryption, are commonly used to ensure that video data remains secure during transmission and processing. User privacy is another critical concern in video anomaly detection systems. Metrics such as the privacy leakage rate evaluate how much sensitive information (e.g., identities, personal activities) can be inferred from the system’s outputs. Differential privacy, federated learning, and encrypted video coding are among the techniques employed to minimize privacy risks. Privacy-preserving methods are evaluated based on the degree of anonymization they provide, often measured in terms of the privacy budget, which balances privacy protection against utility loss.
6.2.3. Robustness-oriented metrics
Robustness metrics aim to assess the system’s ability to maintain reliable performance in the presence of adversarial conditions, such as network disruptions, data corruption, or malicious attacks. One critical metric is fault tolerance, which measures the system’s capacity to continue operating when certain components fail. This is especially important in distributed settings where failures in edge devices or communication links can affect the overall detection pipeline. Techniques such as redundancy, dynamic task migration, and edge-cloud coordination can enhance the system’s fault tolerance. Adversarial robustness evaluates the system’s resilience to attacks designed to manipulate or mislead the anomaly detection model. Adversarial attacks may involve injecting malicious data, such as perturbing video frames or manipulating model parameters. The robustness against such attacks is typically quantified by the system’s ability to maintain high detection accuracy even when exposed to adversarial perturbations. Metrics such as adversarial success rate and robust accuracy (accuracy under attack) are often used to evaluate this aspect. Lastly, scalability measures the system’s ability to handle increasing workloads, including the number of video streams and distributed clients. This is typically evaluated through stress testing, where system performance is analyzed under different load conditions to ensure it can maintain efficiency and reliability as deployment scales up. A scalable NSVAD system should efficiently distribute workloads across edge devices and cloud servers without degrading performance.
@@ 7. Research Cases
To intuitively demonstrate the deployment potential and challenges faced by NSVAD in real-world applications, we selected three of our research cases to apply NSVAD in modern industry (Liu et al., 2024b), smart cities (Liu et al., 2023e), and large-scale systems.
@@ 7.1. Appearance-Motion Prototype Network for Smart Industry
NSVAD can capture global spatial-temporal information in open spaces and detect various anomalous events in industrial settings to ensure production safety and stability. Compared to traditional monitoring devices, video camera data provides stable quality and high information density, making NSVAD the preferred solution for Automatic Industrial Anomaly Detection Systems (AIADS). However, current VAD research lacks a deep focus on industrial scenarios, and existing methods often target outdoor public spaces, struggling with the high-temporal dynamics and complex spatial-temporal interactions in manufacturing environments. The scarcity of industrial video datasets limits AIADS development, while the diversity and complexity of industrial events make it challenging to develop effective NSVAD systems capable of representing regular patterns and generalizing to anomalies. Inspired by the UVAD anomaly categorization in Section 2.2 and the multi-stream approach in Section 3.1, we propose the first NSVAD model for AIADS, named AMP-Net, as shown in Fig. 9 (Liu et al., 2024b). Experiments on real-world industrial datasets show AMP-Net’s ability to handle dynamic environments while maintaining limited generalization to industrial anomalies.
As the first NSVAD model for industrial scenarios, AMP-Net demonstrates the feasibility of using video for automated production monitoring and the potential of NSVAD in smart industries. Its development was influenced by earlier work on multi-stream approaches (Liu et al., 2022c) and memory networks (Gong et al., 2019; Park et al., 2020), which will inspire future researchers designing NSVAD systems for specific applications, as explained in Section 3.1 of Supplement. Although AMP-Net performs well on self-collected industrial datasets, it does not account for deployment costs on terminal devices or handle arbitrary data shifts in real-world environments, indicating further research is needed before it can be widely applied in modern factories. Additionally, analyzing the model’s detection capabilities for different industrial anomaly events and improving its detection of highly hazardous anomalies remain important goals.
@@ 7.2. Causal Video Normality Learning Towards Smart Cities
Smart cities are reshaping lifestyles through AI and next-generation communication, relying on data sensing and understanding. Among the multimodal data from applications like modern transportation, smart homes, and public services, videos dominate in volume and information density, making automated video processing fundamental to smart city research. NSVAD, designed for smart cities, promises solutions for traffic accident detection (Liu et al., 2023c), elderly monitoring (Li et al., 2024b), and public safety. However, challenges like label-independent data shifts and subtle anomalies from event diversity remain. UVAD methods (Liu et al., 2018b; Gong et al., 2019; Chang et al., 2020), including AMP-Net (Liu et al., 2024b), have aimed to balance representation and generalization, but deep learning models still struggle to establish simple mappings between regular events and labels, lacking exploration of causal relationships in video regularity. Our case studies (Liu et al., 2023e) show that deep representation learning (DRL) often fails to distinguish normal and abnormal events, leading to performance degradation in complex scenarios.
To address this, we introduced Causal Representation Learning (CRL) into NSVAD, proposing Causal Video Normality Learning (CVNL) to explore causal relationships (Liu et al., 2023e). Unlike DRL’s statistical dependency-based description, CVNL identifies anomalies based on causal consistency in regular events, enabling better generalization to diverse regular events while remaining sensitive to anomalies. Extensive experiments on large-scale cross-scene datasets show that CVNL meets detection requirements in complex environments like smart cities, maintaining stable performance even with limited regular event data. The simplified CVNL framework is shown in Fig. 10, with more details in (Liu et al., 2023e) and Section 3.2 of Supplement.
@@ 7.3. End-to-Cloud Collaborative Framework for NSVAD
Inspired by NSAI (Song et al., 2022), we propose a hierarchical NSVAD framework for large-scale IoT and mobile internet applications, including hardware, system, algorithm, and application layers. Current NSVAD research focuses mainly on algorithms and applications, such as unsupervised crowd anomaly detection and multimodal violence detection. However, limited computational resources on terminal cameras restrict anomaly detection to short-term, localized events. For example, smart city traffic management requires large-scale systems to monitor traffic flows across a metropolitan area, but current methods only handle local detection at intersections or single devices.
While mobile communication allows devices like smartphones to use cloud servers for large-scale detection, raw data uploads significantly increase communication costs and raise privacy concerns. To address these challenges, we design a configurable end-to-cloud collaborative framework, balancing communication overhead and enhancing data security.
This framework enables terminal devices and cloud servers to collaborate on detecting various anomalies, including local, global, short-term, and long-term events. End devices, such as cameras or mobile devices, handle preprocessing and feature extraction, transmitting only feature data to the cloud, reducing sensitivity and privacy risks. The cloud aggregates features from multiple devices to build large models capable of global, long-term detection and offloads models or results back to the end devices as needed. The framework is flexible, allowing an unlimited number of devices to join and select execution tasks based on real-time performance and complexity, minimizing latency and communication costs while leveraging cloud computing for accurate detection without compromising data privacy.
@@
Key components include computing containers, a deployment platform, and a data pipeline. The containers provide cross-platform, lightweight execution environments for NSVAD models with continuous updates. The deployment platform targets terminal devices like cameras and smartphones, locally deploying large models trained in the cloud. To optimize data transmission, we propose a lightweight data pipeline inspired by Alibaba’s Walle framework, reducing redundancy and communication overhead. The real-time channel uses optimized SSL and long-link communication to minimize latency. Case studies show that our framework reduces cloud workload by 78%, improves detection accuracy (frame-level AUC) by 12%, and lowers false positive rates by 46% compared to traditional cloud-based solutions.
@@ 8.1. Research Challenges
Sections 3-5 introduced the key challenges addressed by various NSVAD algorithms. Here, we further elaborate on unresolved problems from the perspectives of data, labels, models, and systems. In contrast to previous works that emphasize only algorithm design, we also explore the NSVAD-specialized bottlenecks encountered in real-world deployments, such as communication and computing overhead, large-scale detection demand, and privacy concerns.
@@ 8.1.1. Data
Real-world videos exhibit label-independent domain shifts due to variations in scenes, camera angles, and device configurations (Duong et al., 2023). These subtle differences in spatial-temporal patterns, while easily comprehended by humans, often lead to high false positive rates in NSVAD models (Liu et al., 2023e). Existing methods typically validate their models on datasets from a single scene to avoid this issue (Ramachandra et al., 2020). For instance, the UCSD dataset (Li et al., 2013) includes two distinct perspectives, but they are treated as separate datasets. The ShanghaiTech dataset (Liu et al., 2018b), despite spanning 13 scenes, is often treated as a single scene, leading to performance drops when compared to simpler datasets like UCSD Ped2 (Li et al., 2013) and CUHK Avenue (Lu et al., 2013). In real-world applications, such scene and device variations are inevitable, making it impractical to develop specialized models for every setup. To address this, some researchers (Acsintoae et al., 2022) have proposed using virtual engines to simulate anomaly events and generate richer positive samples. However, datasets like XD-Violence (Wu et al., 2020), which include movie and game scenes, differ from real-world anomalies, limiting the effectiveness of models trained on them. Bridging this domain gap between virtual and real anomalies is essential for deployable NSVAD systems. Additionally, multimodal NSVAD, while a growing field, remains confined to the fusion of RGB images and synchronized audio, neglecting novel modalities like language (Wu et al., 2024b) and texts (Pu et al., 2024), limiting its application in media streaming and broadcasting.
@@ 8.1.2. Label
Despite the acknowledged rarity and diversity of anomalies, most models are trained unsupervised (Yang et al., 2023a; Yan et al., 2023; Wang et al., 2023), using only normal samples. However, as discussed in Section 2.1, unsupervised models still require anomaly-free samples during training. Challenges include the impact of pixel noise on model performance and the cost of labeling large-scale datasets for supervised methods. Moreover, UVAD methods rely on data filtering to avoid contaminated training sets, preventing online learning directly from raw video streams, as seen in FuVAD (Pang et al., 2020). Conversely, WsVAD utilizes video-level labels to reduce labor costs (Sultani et al., 2018), but improving the stability of FuVAD/UVAD in complex environments or creating hybrid models that can mitigate label noise remain future directions.
@@ 8.1.3. Model
Unsupervised NSVAD models benefit from self-supervised learning for spatial-temporal feature representation but face challenges with over-generalization, which can lead to missed anomalies (Park et al., 2020). The key challenge is balancing representation and generalization to reduce both false positives and negatives (Liu et al., 2024b). Approaches like memory networks and causal representation learning (Liu et al., 2023e) have shown promise, but performance on complex datasets remains inconsistent. In contrast, WsVAD can only detect predefined anomaly events, limiting its adaptability and failing to meet open-world requirements. While WsVAD models generally offer more reliable results than UVAD and FuVAD, their reliance on ranking loss is still debated. To advance NSVAD, there is a need for interpretable models and systems that integrate data encryption to address privacy concerns.
@@ 8.1.4. System
Current NSVAD research primarily focuses on improving algorithmic performance on existing datasets, overlooking practical deployment challenges like device heterogeneity and resource constraints. Different applications, such as surveillance analysis, content monitoring, and non-real-time video detection, require distinct task setups and data modalities. However, most work has concentrated on surveillance video detection, neglecting applications like short video analysis and live streaming. Additionally, mobile devices, with their limited storage, computing, and communication resources, necessitate lightweight model designs. Moving forward, NSVAD should focus on the synchronous optimization of detection performance and system cost.
8.1.5. NSVAD-specialized
For real-world applications, NSVAD needs to handle vast data streams from urban surveillance cameras while addressing long-term anomaly detection. These systems also face public concerns regarding data overhead and privacy security. Unlike algorithmic research, NSVAD for large-scale applications presents three distinct challenges: (i) Efficient data exchange and task offloading (Dong et al., 2024) between millions of cameras and distributed servers. Current research focuses on enhancing algorithms but neglects the costs of data acquisition and transmission, which are critical in large-scale deployments. NSVAD must adopt new video compression (Gomes et al., 2023) and transmission protocols to handle the expanding network of cameras and increasing video resolution. Distributed machine learning techniques can also facilitate global model training without aggregating all data. (ii) Detecting anomalies across large spatial and temporal scales. Video IoT systems collect information from entire cities, but current NSVAD models (Su et al., 2025; Zhou et al., 2024), designed for discrete scenes, struggle with large-scale data. For example, traffic anomalies like congestion may be easily detected at intersections, but large-scale crowd movements might be misinterpreted as anomalous gatherings. (iii) Privacy protection and ethical concerns. NSVAD systems collect identifiable information such as faces and clothing, raising concerns about privacy breaches and biases related to race, gender, and skin color (Su et al., 2023, 2023; Noghre et al., 2024). While identity-agnostic data (e.g., encrypted video (Cheng et al., 2020)) and privacy-preserving techniques (e.g., federated learning (Al-Lahham et al., 2024)) are promising, ensuring transparency and fairness remains a significant challenge for NSVAD development.